Graph the open sentence. I need help with this. Thank you.
Determine an equation in the form y=mx+b of the line passing through (-2, 5) and parallel to the line 2x+y=6.

Answer by Alan3354(69443)

Do it like this one:
----
How would u find a line that is parallel to y=(2/3)x+5, passing through (4,-5)
----------
Since the lines are parallel, the slopes are equal, = 2/3
------
y+5 = (2/3)*(x-4)
3y + 15 = 2x - 8
2x - 3y = 23